# Whom We Trust, What We Fear: COVID-19 Fear and the Politics of Information

### v1

@@ 4. Statistical analysis of fear
Figure 2. Temporal evolution of the singleton fear scores $\phi_{s^{\prime}}$—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. To reduce noise, official infections and $\phi_{s^{\prime}}$ were smoothed using a two-month rolling average.
@@
Employing the definitions of the previous section, we investigate the drivers behind the evolution of the nationwide U.S. fear score $\phi_{s^{\prime}}$. As a starting point, we examine whether Fear differs across the various information sources. In particular, considering the nine distributions derived from the singleton information sources $s^{\prime}$, we first applied the Shapiro–Wilk test (Shapiro and Wilk, 1965) to assess the normality of each distribution, which indicated that none of the distributions follow a normal distribution. Subsequently, to evaluate whether these distributions differ significantly from one another—including both intra-group comparisons (i.e., across the nine singleton information sources within a given Age or Education group, as well as at the population level) and inter-group comparisons (i.e., comparing the nine singleton information source across different Age or Education groups)—, we performed the Kolmogorov–Smirnov test (Massey Jr, 1951)—it considers differences across the entire distribution—and the Mann–Whitney U test (Nachar et al., 2008)—which is primarily sensitive to differences in central tendency. The consistent significance across comparisons suggests meaningful variation in fear score distributions across both intra-group and inter-group levels, underscoring the influence of demographic and informational heterogeneity.
Figure 3. Spearman correlation coefficients among the time series shown in Figure 2.
@@
Some sources show even greater sensitivity. The coefficient of variation [^3] The coefficient of variation is defined as the standard deviation divided by the absolute value of the mean, providing a measure of relative variability. is particularly high for Scientists ($\approx 3$) and the CDC ($\approx 1$), whereas all other sources remain below $0.5$, indicating moderate variability. This pattern illustrates how fear intensifies with the rising visibility of COVID-19 cases, dominating public discourse. The moderate to high variation highlights its strong dependence on fluctuations in official case numbers.
Among all sources, it is noteworthy that Family and friends and Journalists slightly anticipate phase transitions in the evolution of infections: the signal $\phi_{s^{\prime}}$ precedes both the initial rise in case numbers during the outbreak and the subsequent decline at the peak for each of the two major waves—corresponding to the Delta and Omicron variant surges (Dattani and Rodés‑Guirao, 2024). This pattern may reflect an anticipatory response based on early information or heightened awareness within these social and expert networks.
In addition, singleton fear scores exhibit strong mutual correlations, as demonstrated by the Spearman correlation coefficients (Spearman, 1961) in Figure 3. A strong correlation among the fear-related time series indicates that, regardless of the specific source of information, the perceived fear tends to vary consistently across sources. A lower correlation is observed for the two outliers mentioned above, as well as for None of the above, which exhibits a flatter trend. However, these three time series are highly correlated with one another and with Journalists, indicating that individuals relying on personal or alternative information sources may develop perceptions of fear that differ from those shaped by institutional or official sources. This may point to the existence of an alternative information ecosystem—insulated from broader media trends—where emotional responses are shaped by interpersonal communication or internal beliefs rather than external news flows.
Figure 4. Boxplots showing the distribution of the average fear score across the nine singleton information sources, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups.
Finally, Figure 4 displays boxplots illustrating the distribution of the average fear score across the nine singleton information sources, stratified by Age and Education groups. By excluding the temporal dimension, these distributions offer a concise statistical summary of how fear scores vary across different age and education strata. Additionally, it includes the overall distribution without stratification by Age or Education—note that the overall boxplot is not merely the average of the stratified ones, as the estimates of $\phi_{s^{\prime}}$ were obtained using separate linear models for each grouping variable, as detailed in Equation 5. Specifically, focusing on Age, the average fear score increases with age, with $Age_{1}$ exhibiting the lowest fear score and $Age_{3}$ the highest, across almost all sources—a few exceptions to this trend are observed for CDC, Doctors, and Scientists, where the pattern differs slightly. When focusing on Education, the average fear score generally decreases from $Edu_{1}$ to $Edu_{2}$. In most cases, the fear score then increases again with $Edu_{3}$, although there are some exceptions where it continues to decrease. This behavior highlights the nuanced effect of education on perceived fear, suggesting that while higher education may generally be associated with greater awareness and concern, its influence can vary depending on the source of information and potentially other demographic or contextual factors. In addition, Religious leaders and Family and friends continue to stand out as outliers in the distribution.
5. Causal inference model
The findings from the previous section indicate that the fear score is strongly correlated not only with the sources from which people obtain their information but also with their age and education. Together, these factors contribute significantly to variations in fear across demographic groups. Nonetheless, identifying the primary driver of perceived fear remains challenging. Therefore, we employ causal inference methods to investigate these factors and determine which variable among Source, Age, and Education has the strongest influence on Fear.
Specifically, causal inference involves identifying cause-and-effect relationships by distinguishing them from correlations. It employs statistical methods and study designs—such as randomized controlled trials and quasi-experimental approaches—to estimate the impact of an exposure or intervention on an outcome, while also accounting for confounding factors. While randomized controlled trials represent the gold standard for establishing causality, they are often impractical or unethical. In such cases, causal inference methods can be applied to observational data, provided certain conditions are met—for example, that all relevant variables influencing both the treatment and the outcome are observed, and that the effect of the treatment is consistent across individuals. Common approaches include propensity score matching, instrumental variable analysis, and structural causal models (Pearl, 2010).
Figure 5. Causal inference model used.
Figure 5 illustrates the causal inference model employed in this study, involving four variables: Fear, Source, Age, and Education. Specifically: i) Source $\rightarrow$ Fear denotes the direct causal effect of information source on the outcome variable, the fear score; ii) Age $\rightarrow$ Fear and Education $\rightarrow$ Fear indicate a direct causal pathway from demographic factors to Fear, capturing age- and education-related differences in Fear independently of other influences; iii) Age $\rightarrow$ Source and Education $\rightarrow$ Source represent confounding paths where demographic factors influence the distribution or assignment of individuals to different information sources, potentially confounding the effect of Source on Fear; and iv) Age $\rightarrow$ Education reflects the causal effect of Age on Education level. Age, Education, and Source are categorical variables with Age, Education $\in$ {1, 2, 3} and Source $\in$ {1, 2, …, 9}—each representing one of the nine singleton information sources—, while Fear is a continuous variable.
To define and analyze the causal inference model, we employed the DoWhy package (Sharma and Kiciman, 2020; Blöbaum et al., 2024), a Python library that combines graphical models and statistical methods to specify assumptions, identify causal effects, and estimate them using various techniques. It also provides tools for robustness testing and counterfactual analysis, enabling rigorous causal inference from observational data.
We systematically estimate the strength of each direct causal pathway leading to Fear. This approach allows us to quantify the individual contribution of each variable to variations in Fear, providing a clearer understanding of the underlying causal structure. Our analysis reveals that Source emerges as the most influential variable, explaining a substantial 73.8% of the variance in Fear. This indicates that this variable plays a dominant role in shaping fear responses within the population studied. In comparison, demographic factors such as Age and Education contribute more modestly, accounting for 6.9% and 6.6% of the variance, respectively. These findings suggest that while individual characteristics do have some impact, the Source variable is the primary driver. The remaining 12.7% of variance is attributed to other factors or unobserved variables not included in our current model, highlighting the potential influence of additional elements beyond those considered. This comprehensive estimation underscores the value of causal inference methods in disentangling complex relationships within observational data.
6. Geographical distribution of sources
Figure 6. Temporal evolution of the singleton information sources $\hat{\sigma}_{s^{\prime}}$—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
U.S. nationwide. Using the definitions introduced in Section 3, we analyze patterns in the use of each singleton information source, independently of Fear, represented by $\hat{\sigma}_{s^{\prime}}$. As before, we begin by analyzing the nine distributions; the Shapiro–Wilk test indicates that none follow a normal distribution. To assess differences—including both intra-group and inter-group comparisons—we used the Kolmogorov–Smirnov and Mann–Whitney U tests. The observed significant differences reinforce the idea that perceived uncertainty varies meaningfully across demographic segments and source groupings.
Based on this evidence, Figure 6 shows the evolution of the fraction of respondents relying on each information source, alongside the official number of infections across the U.S. It is notable that None of the above consistently shows values near zero with minimal variation, appearing as an outlier; therefore, we chose not to include it in the visualization. This suggests that very few survey respondents rely on information sources outside those listed in the question. Similarly to the fear score, the reliance of respondents on information sources closely tracks infection trends, exhibiting strong correlation—as shown in Figure 7—and increased sensitivity to relative increases in case numbers rather than absolute values—with the notable exception of Politicians, whose pattern remains flatter and largely uncorrelated. Some sources exhibit moderate sensitivity. The coefficient of variation varies between approximately $0.25$ for None of the above to around $0.1$ for Doctors and CDC, and is lower for other sources, indicating moderate to low variability.
@@ 6. Geographical distribution of sources
Figure 7. Spearman correlation coefficients among the time series shown in Figure 6.
@@
Figure 8. Boxplots showing the distribution of the fraction of respondents relying on each information source, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
@@
Finally, Figure 8 displays boxplots illustrating the distribution of the fraction of individuals relying on each singleton information source, stratified by Age and Education groups, alongside the overall distribution aggregated without stratification. In this case as well, we chose not to include None of the above in the visualization. In both instances, it stands out as a clear outlier. Three clear patterns emerge when moving from $Age_{1}$ to $Age_{3}$ and from $Edu_{1}$ to $Edu_{3}$. First, certain sources maintain a relatively stable or flat level of reliance across these groups, suggesting that demographic factors have little effect on their usage—like Family and friends, Politicians, and None of the above. Second, Religious leaders exhibit a decreasing trend, with younger or less educated groups relying on them more than older or more educated ones. Third, the remaining sources—institutional or expert ones—show increased reliance as Age and Education rise—meaning older or more educated individuals tend to trust these sources more.
@@
U.S. states. To further understand patterns in the use of information sources, we conduct a geographical analysis across U.S. states. The United States presents a diverse socio-economic landscape, with substantial variation in income levels, educational attainment, internet access, and cultural norms across regions. These differences are likely to influence how individuals seek and trust information, particularly in the context of a public health crisis such as the COVID-19 pandemic. By examining the spatial distribution of preferred information sources, we aim to uncover regional trends and disparities that may reflect underlying socio-demographic and economic characteristics. Hence, for each U.S. state, we compute the average reliance on each singleton information source—from $\hat{\sigma}_{s^{\prime}}$—, without stratification by Age or Education, resulting in nine values per state. These averages are then used to cluster the states using K-means (Hartigan et al., 1979), a fast and widely used clustering algorithm that partitions data into a predefined number of clusters by minimizing the distance to cluster centers—it assumes spherical, equally sized clusters. Clustering quality is often evaluated using the Silhouette score (Shahapure and Nicholas, 2020), which measures how well points fit within their assigned clusters, with higher values indicating better-defined clusters.
@@
Figure 9. Clusters derived from K-means analysis based on average reliance on each information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, without stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections. Misclassified states are overlaid with a contrasting grid. States with a purple border indicate swing states between the 2020 and 2024 U.S. political elections.
@@
Based on the Silhouette score, we determined that two clusters provide the optimal partitioning. Figure 9 illustrates the resulting clusters, with each state colored according to its assigned cluster. Interestingly, by chance, we obtained two clusters that correspond closely to the two major U.S. political affiliations: Democratic and Republican. Specifically, Figure 9.a) compares the derived clusters to the 2020 U.S. political elections results, while Figure 9.b) shows the comparison with the 2024 elections outcomes. States that were misclassified—i.e., assigned to the cluster contrary to their political affiliation in the given year—are marked with a grid overlay in the contrasting cluster color. We observe that only two states—along with Alaska, which is not shown on the map—were assigned to the opposite cluster in the 2020 elections, while four states—again including Alaska—were assigned to the opposite cluster in the 2024 elections. Interestingly, these states—except for Alaska—are swing states between the 2020 and 2024 political elections, which makes them inherently more difficult to classify. These results suggest that the clustering based on information source reliance closely reflects established political divisions, with misclassifications primarily occurring in politically competitive swing states. This underscores the sensitivity of information consumption patterns to shifting political landscapes.
7. Conclusion

### v2

@@ Results
Figure 3: Temporal evolution of $\phi_{s^{\prime}}$ for each of the nine singleton information sources—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. To reduce noise, official infections and $\phi_{s^{\prime}}$ were smoothed using a two-month rolling average.
@@
Statistical analysis on fear. Employing the definitions introduced in the Data section, we investigate the drivers behind the evolution of the nationwide U.S. fear score $\phi_{s^{\prime}}$. As a starting point, we examine whether Fear differs across the various information sources. In particular, considering the nine distributions derived from the singleton information sources $s^{\prime}$, we first applied the Shapiro–Wilk test, which indicated that none of the distributions are normally distributed. Next, to assess whether these distributions differ significantly—both within groups (across the nine singleton information sources for each Age or Education group and at the population level) and between groups (across different Age or Education groups)—we applied the Kolmogorov–Smirnov and Mann–Whitney U tests. The consistently significant results indicate meaningful variation in fear-score distributions at both levels, reflecting demographic and informational heterogeneity.
@@
Figure 4: Spearman correlation coefficients among the time series shown in Figure 3.
Some sources show even greater sensitivity. The coefficient of variation [^3] The coefficient of variation is the standard deviation divided by the mean’s absolute value, measuring relative variability. is particularly high for Scientists ($\approx 3$) and the CDC ($\approx 1$), whereas all other sources remain below $0.5$, indicating moderate variability. This pattern illustrates how fear intensifies with the rising visibility of COVID-19 cases, dominating public discourse. The moderate to high variation highlights its strong dependence on fluctuations in official case numbers.
Among all sources, Family and friends and Journalists slightly anticipate phase transitions in infection trends: their signal $\phi_{s^{\prime}}$ precedes both the initial rise and the peak decline during the Delta and Omicron waves (Dattani and Rodés‑Guirao 2024). This pattern may reflect an anticipatory response driven by early information or heightened awareness within these social and media networks.
In addition, singleton fear scores exhibit strong mutual correlations, as demonstrated by the Spearman correlation coefficients in Figure 4. A strong correlation among the fear-related time series indicates that, regardless of the specific source of information, the perceived fear tends to vary consistently across sources. A lower correlation is observed for the two outliers mentioned above, as well as for None of the above, which exhibits a flatter trend. However, these three time series are highly correlated with one another and with Journalists, indicating that individuals relying on personal or alternative information sources may develop perceptions of fear that differ from those shaped by institutional or official sources. This may point to the existence of an alternative information ecosystem—insulated from broader media trends—where emotional responses are shaped by interpersonal communication or internal beliefs rather than external news flows.
Figure 5: Boxplots showing the distribution of $\phi_{s^{\prime}}$ across the nine singleton information sources, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups.
Finally, Figure 5 displays boxplots illustrating the distribution of the signal $\phi_{s^{\prime}}$ across the nine singleton information sources, stratified by Age and Education groups. By excluding the temporal dimension, these distributions offer a concise statistical summary of how fear scores vary across different age and education strata. Additionally, it includes the overall distribution without stratification by Age or Education—note that the overall boxplot is not merely the average of the stratified ones, as the estimates of $\phi_{s^{\prime}}$ were obtained using separate linear models for each grouping variable, as detailed in Equation 5. Specifically, focusing on Age, $\phi_{s^{\prime}}$ increases with age, with $Age_{1}$ exhibiting the lowest fear score and $Age_{3}$ the highest, across almost all sources—a few exceptions to this trend are observed for CDC, Doctors, and Scientists, where the pattern differs slightly. When focusing on Education, $\phi_{s^{\prime}}$ generally decreases from $Edu_{1}$ to $Edu_{2}$, and in most cases rises again with $Edu_{3}$, although in some instances it continues to decline. This behavior highlights the nuanced effect of education on perceived fear, suggesting that while higher education may generally be associated with greater awareness and concern, its influence can vary depending on the source of information and potentially other demographic or contextual factors. In addition, Religious leaders and Family and friends continue to stand out as outliers in the distribution.
Causal inference. The findings from the previous section indicate that $\phi_{s^{\prime}}$ varies significantly not only with the sources of information people rely on but also with their age and education. Nonetheless, identifying the primary driver of perceived fear remains challenging. Therefore, we employ causal inference methods to investigate these factors and determine which variable among Source, Age, and Education has the strongest influence on Fear.
Figure 6: Estimated direct effects on Fear. Panels a) and b) show variance explained by each variable in the first and second waves, respectively. Panels c) and d) show the influence of each information source.
Figure 2 illustrates the causal inference model, featuring nine binary Source variables (one-hot encoded from the original categorical source), continuous outcome Fear, and categorical confounders Age, Education $\in\{1,2,3\}$. Direct effects flow from each Source $\rightarrow$ Fear and demographics $\textit{Age}{},\textit{Education}{}\rightarrow$ Fear. Critical confounding arises via $\textit{Age}{},\textit{Education}{}\rightarrow$ Source, where demographics determine information source exposure, biasing naive source-fear associations; additionally, $\textit{Age}{}\rightarrow$ Education captures age’s effect on education. This DAG structure necessitates adjustment for both confounders to identify the causal effects of information sources on fear. We exclude infection counts in the model because we assume that individuals learn about epidemiological conditions exclusively through information sources. In the absence of such sources, individuals would have no access to the actual infection counts. To account for temporal dynamics, we analyze two distinct waves, separated by the date on which infections reached their minimum between the two periods under consideration: May 2021–November 11, 2021, and November 12, 2021–June 2022. Within each wave-specific analysis, we acknowledge that daily observations may exhibit temporal correlation.
Source ATE Source ATE
W1 W2 W1 W2
Scientists 0.24 0.37 Doctors -0.17 0.00
Government 0.18 0.25 Journalists -0.38 -0.25
CDC 0.18 0.33 Religious leaders -1.62 -1.50
Politicians 0.14 0.24 Family and friends -1.90 -1.89
Table 3: ATE of information sources on Fear across two waves using None of the above as the “control” source. W1 and W2 denote the first and second waves of the data.
Specifically, we applied this model—normalizing $\phi_{s^{\prime}}$—to estimate the strength of each direct causal pathway leading to Fear. Our analysis shows that Source is the most influential variable, accounting for 81.8% of the normalized $\phi_{s^{\prime}}$ variance in the first wave and 81.5% in the second—Figures 6.a) and 6.b). This indicates that this variable plays a dominant role in shaping fear responses within the population studied. In comparison, demographic factors such as Age and Education contribute more modestly, accounting for 6.7%-8.8% and 11.5%-9.7% of the variance, respectively, across the two waves. These findings suggest that while individual characteristics do have some impact, the Source variable is the primary driver. Moreover, Figures 6.c) and 6.d) show the influence of each Source, with Family and friends and Religious leaders exerting the greatest influence, while all others have smaller effects—except for Doctors in the first wave, which account for 6.3%.
Finally, we estimate the ATE of different information sources on Fear, using None of the above as the “control” source. Table 3 presents the results. In both waves, the ranking of sources in decreasing order is largely consistent, with the exception of CDC in the second wave, whose effect is higher and closer to that of Scientists. Interestingly, from the first to the second wave, the ATE of each source increases, indicating that during the Omicron wave, people were more afraid when using a source other than the “control” source. Notably, we observe that people are less afraid when obtaining information from the two outliers, Family and friends and Religious leaders, as well as from Journalists, suggesting lower trust in these sources, whereas they are more afraid when obtaining information from more official sources such as Scientists, Government, CDC, and Politicians.
Geographical distribution of sources: U.S. nationwide. Using the definitions introduced in the Data section, we analyze patterns in the use of each singleton information source, independently of Fear, represented by $\hat{\sigma}_{s^{\prime}}$. As before, we begin by analyzing the nine distributions: the Shapiro–Wilk test indicates that none follow a normal distribution. To assess differences—both within and between groups—we used the Kolmogorov–Smirnov and Mann–Whitney U tests. The observed significant differences reinforce the idea that perceived uncertainty varies meaningfully across demographic segments and source groupings.
Based on this evidence, Figure 7 shows the evolution of the fraction of respondents relying on each information source, alongside the official number of infections across the U.S. It is notable that None of the above consistently shows values near zero with little variation and is therefore excluded from the visualization, indicating that few respondents rely on sources outside those listed. Similar to the fear score, respondents’ reliance on information sources closely follows infection trends, showing strong correlation and greater sensitivity to relative increases in cases, except for Politicians, whose pattern remains largely flat and uncorrelated—Figure 8. Some sources exhibit moderate sensitivity. The coefficient of variation varies between approximately $0.25$ for None of the above to around $0.1$ for Doctors and CDC, and is lower for other sources, indicating moderate to low variability.
Figure 7: Temporal evolution of $\hat{\sigma}_{s^{\prime}}$ for each of the nine singleton information sources—values on the left axis—, alongside the official infection counts in the U.S.—values on the right axis—between May 2021 and June 2022. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
@@
Figure 8: Spearman correlation coefficients among the time series shown in Figure 7.
@@
Finally, Figure 9 displays boxplots illustrating the distribution of the fraction of individuals relying on each singleton information source, stratified by Age and Education groups, alongside the overall distribution aggregated without stratification. In this case as well, we excluded None of the above from the visualization, as it stands out as a clear outlier in both instances. Three clear patterns emerge when moving from $Age_{1}$ to $Age_{3}$ and from $Edu_{1}$ to $Edu_{3}$. First, certain sources maintain a relatively stable or flat level of reliance across these groups, suggesting that demographic factors have little effect on their usage—like Family and friends, Politicians, and None of the above. Second, Religious leaders exhibit a decreasing trend, with younger or less educated groups relying on them more than older or more educated ones. Third, the remaining sources—institutional or expert ones—show increased reliance as Age and Education rise—meaning older or more educated individuals tend to trust these sources more.
Figure 9: Boxplots showing the distribution of the fraction of respondents relying on each information source, stratified by a) Age—where $Age_{1}$, $Age_{2}$, and $Age_{3}$ refers to the three Age groups—and b) Education groups—where $Edu_{1}$, $Edu_{2}$, and $Edu_{3}$ refers to the three Education groups. We chose not to display None of the above as its values are consistently near zero and exhibit minimal variation.
@@
Geographical distribution of sources: U.S. states. To further understand patterns in the use of information sources, we conduct a geographical analysis across U.S. states. The United States presents a diverse socio-economic landscape, with substantial variation in income levels, educational attainment, and cultural norms across regions. These differences are likely to influence how individuals seek and trust information, particularly in the context of a public health crisis such as the COVID-19 pandemic. By examining the spatial distribution of preferred information sources, we aim to uncover regional trends and disparities that may reflect underlying socio-demographic and economic characteristics. Hence, for each U.S. state, we compute the average reliance on each singleton information source—from $\hat{\sigma}_{s^{\prime}}$—, without stratification by Age or Education, resulting in nine values per state. These averages are then used to cluster the states using K-means, spectral, and hierarchical clustering methods.
@@
Figure 10: Clusters derived from K-means analysis based on average reliance on each information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, without stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections. Misclassified states are overlaid with a contrasting grid. States with a purple border indicate swing states between the 2020 and 2024 U.S. political elections.
@@
Based on the Silhouette score, we determined that two clusters provide the optimal partitioning. Figure 10 illustrates the resulting clusters, with each state colored according to its assigned cluster. Interestingly, by chance, we obtained two clusters that correspond closely to the two major U.S. political affiliations: Democratic and Republican. Specifically, Figure 10.a) compares the derived clusters to the 2020 U.S. political elections results, while Figure 10.b) shows the comparison with the 2024 elections outcomes, both obtained using K-means clustering—similar results were obtained with hierarchical and spectral clustering. States that were misclassified—i.e., assigned to the cluster contrary to their political affiliation in the given year—are marked with a grid overlay in the contrasting cluster color. We observe that only three states were assigned to the opposite cluster in the 2020 elections, while five states were assigned to the opposite cluster in the 2024 elections. Interestingly, these states—except for Alaska—are swing states between the 2020 and 2024 political elections, which makes them inherently more difficult to classify. These results suggest that the clustering based on information source reliance closely reflects established political divisions, with misclassifications primarily occurring in politically competitive swing states. This underscores the sensitivity of information consumption patterns to shifting political landscapes.
Conclusion